# Triples as Links between Binary Black Hole Mergers, their Electromagnetic Counterparts, and Galactic Black Holes

### v1

@@ 1 Introduction
Stellar-mass black holes (BHs) can be detected through multiple observational methods, each of which potentially probes different evolutionary pathways. The oldest method relies on electromagnetic (EM) signatures, most notably X-ray emission from accretion disks in X-ray binaries. These systems, discovered in the 1960s, provided the first indirect evidence for stellar-mass BHs (e.g., Bolton, 1972; Remillard & McClintock, 2006; MacLeod & Grindlay, 2023). More recently, the detection of gravitational waves (GWs) from merging binary BHs by the LIGO–Virgo–KAGRA (LVK) Collaboration has revolutionized our understanding of the BH mass distribution (e.g., Abbott et al., 2016), revealing a population of heavy BHs (up to $\sim 100$ M) without any known EM activity [^1] A handful of LVK BHs have been suggested to coincide with EM flares in AGN (Graham et al., 2023) but their brightness is difficult to explain (Tagawa et al., 2024) and the significance of the spatial associations are controversial (Veronesi et al., 2025).. The third avenue of detection, based on astrometric measurements of orbital motion, has been enabled by Gaia, which has opened a new observational window into detached BH binaries (El-Badry et al., 2023b). Stellar-mass BHs can also be detected by microlensing (e.g., Agol et al., 2002; Lu et al., 2016; Lam et al., 2022; Sahu et al., 2022) or by combined radial velocity and photometric obeservations of non-accreting BHs in binaries (e.g., Thompson et al., 2019; Liu et al., 2019b; Clavel et al., 2021; Chawla et al., 2024), but our focus in this paper is on the first three channels summarized above.
@@
We stress that while hierarchical triples have been proposed as a formation channel for merging binary BHs in the field, their overall contribution to the LVK-detected merger rate is typically estimated to be small and suppressed by a factor of $\sim 5$$30$ relative to the estimated detected rate (e.g., Antonini et al., 2016, 2017; Silsbee & Tremaine, 2017; Liu et al., 2019a; Kummer et al., 2025). Nonetheless, triple evolution may naturally account for certain systems that are otherwise difficult to explain, such as BH mergers in the mass gap (e.g., Vigna-Gómez et al., 2021a; Martinez et al., 2022). In this work, however, we focus on a complementary scenario: BH binaries that merge through isolated binary evolution, with a tertiary companion that does not play a significant dynamical role in driving the merger. In particular, the tertiary star’s presence is relevant only after the inner binary merges, when the system receives a recoil kick. Our setup, therefore, explores the consequences of a passive tertiary in an otherwise binary-driven channel. We assume throughout this paper that the distant tertiary companion remains bound during the formation of the two BHs in the inner binary, and focus on the dynamics that follow when the inner BH binary merges. The validity of this assumption is sensitive, of course, to the uncertain mass ejection and natal kick during BH formation. Recent detections of X-ray binaries with distant companions imply that at least some distant companions remain bound during BH formation (e.g., Burdge et al., 2024; Shariat et al., 2025b).
@@
Figure 2: Two examples of the orbital outcome of a recoil kick on an orbit. We consider a system of two low-spin BHs with $S_{1}=S_{2}=0.01$, and spin-orbit angles of $2^{\circ}$, $1^{\circ}$, $\Omega_{s1}=90^{\circ}$ and $\Omega_{s2}=270^{\circ}$. The mutual inclination is set to $0.1^{\circ}$. Before the kick, the star has a semi-major axis of $20$ au and eccentricity $e_{\star}=0.9$ (right column), and $e_{\star}=0.83$ (left column). The argument of periapsis of the inner (outer) binary at the time of the merger is $300^{\circ}$ ($20^{\circ}$). These angles are relevant for the rotation of the various vectors to the invariant plane. See text for more details. We assume that the kick took place when the star’s phase was at $10^{\circ}$ (right column), and $5^{\circ}$ (left column). Various post-merger quantities are shown as a function of the mass ratio of the two original BHs, $q$. The bottom row shows the closest approach, Equation (3). Overplotted are the tidal radius, $r_{\rm circ}$, and the Roche radius $r_{\rm Roche}$. We expect EM signatures for BHs that approach their stellar companion within these radii. The middle row depicts the post-kick semi major axis $a_{\star,n}$, solid red lines and pericenter $a_{\star,n}(1-e_{\star,n})$, dashed purple lines. Overplotted are the initial (pre-kick) semi-major axis and pericenter of the star, solid and dashed gray lines, respectively. Lastly, Gaia-BH1’s semi-major axis and pericenter ($a_{\rm Gaia~BH1}$, and $r_{p,\rm Gaia~BH1}=a_{\rm Gaia~BH1}(1-e_{\rm Gaia~BH1})$) are also overplotted, in pink and light purple lines (El-Badry et al., 2023b). The top row shows the normalized velocity $u_{k}=v_{\rm kick}/v_{r}$, and $\sin\gamma$, where $\gamma$ is the angle between the radius vector and the post-kick velocity vector. The grid highlights that when $u_{k}=1$, $\gamma\to 0^{\circ}$, thus $R_{c}\to 0$.
@@ 2.1 The Triple Configuration
Throughout this work, we consider a stellar-mass BH binary, $m_{1}$ and $m_{2}$, where the mass ratio is $q\equiv m_{1}/m_{2}\leq 1$, where “1” stands for the less massive component. The total mass of the binary is $M=m_{1}+m_{2}$, semi-major axis $a_{\rm BHs}$ and eccentricity $e_{\rm BH}$, associated with angular momentum $L_{\rm BHs}$. We define the spin-orbit angle of mass $m_{j}$ as $i_{sj}=\hat{h}\cdot\hat{S}_{j}$, where $\hat{h}$ is the unit vector along the BH’s orbital angular momentum, and $j\in\{1,2\}$. The corresponding spherical coordinate angles are $\Omega_{sj}$ for each mass. Further, this binary is orbited by a star with a mass $m_{\star}=1$ M, and semi-major axis $a_{\star}$ and eccentricity $e_{\star}$. We note that while triples are observed to have a twin excess (Shariat et al., 2025a), the long lifetime of 11 M benefits the calculation. A full IMF calculation is beyond the scope of this paper. The frame of reference considered here is the inertial frame, for which the $z$-axis is parallel to the total angular momentum. After the BHs merge, they typically receive a recoil kick due to the anisotropic emission of GWs.
@@ 2.2 Impact on the Stellar Orbit
We adopt the Lousto et al. (2010, 2012) fitting formulae, which provide the recoil kick velocity. See Appendix A for the relevant set of equations. In this Letter, we focus on the merger channel driven by binary stellar evolution. This channel implies that the BH spins are aligned with its angular momentum. As a result, a well-known outcome emerges: the recoil kick is directed within the plane of the BH binary (e.g., Lousto et al., 2012). We discuss the effects of the spin-alignment approximation in Section 4.3.
@@
(bottom dashed line). A BH approaching the star inside either of these limits will likely result in an EM source. Orbits with $R_{c}\lesssim r_{\rm Roche}$ will result in the tidal disruption of the star (or, in some cases, a direct collision, e.g., Perets et al., 2016; Fragione & Leigh, 2018; Rose et al., 2022; Yang et al., 2022; Ryu et al., 2022, 2023, 2024; Xin et al., 2024). Orbits with $R_{c}\lesssim r_{\rm circ}$ will lead to the orbit circularizing by tides, eventually producing a mass transferring system such as a low-mass X-ray binary. We discuss these outcomes in more detail below. Note that $r_{\rm circ}$ is related to the tidal capture radius Press & Teukolsky (1977), but in our case, most of the post-BH merger orbits are already bound, so there is no “capture.” The relevant radius is instead the pericenter distance inside of which tides lead to circularization on a timescale less than the evolutionary time of the star. This is likely a factor of a few larger than the traditional tidal capture radius used in Equation 8 (see, e.g. Wu, 2018), so our estimates of the fraction of the stars producing X-ray binaries are likely conservative.
@@ 3 A population study
As a proof of concept, we focus on a population of isolated BH binaries that will ultimately become LVK sources, and add a stellar tertiary to them. The isolated binary channel involves a common-envelope, stable mass transfer phase, or chemically homogeneous evolution (e.g., Belczynski et al., 2002, 2007; Dominik et al., 2012, 2015; Stevenson et al., 2017; de Mink & Mandel, 2016; Mandel & de Mink, 2016; Gallegos-Garcia et al., 2021). We thus adopt binary BH orbital configurations that isolated binary population synthesis work suggests will merge. Specifically, we are motivated by the angular momentum constraints from Figure 13 in Kruckow et al. (2018), which represent the orbital configurations of two BHs, right after the formation of the second BH. Specifically, we choose the BH binary eccentricity from a uniform distribution and apply the angular momentum constraints on the binary’s semi-major axis. Additionally, the population is limited to a merger time of $\leq 10$ Gyr (which reduces the population by a factor of 2).
@@
For simplicity, the BH masses are adopted from a uniform distribution, each between 91009-100 M. For the outer orbit, we are motivated by triple conditions that include post-main-sequence stellar evolution for massive stars (e.g., Naoz et al., 2016; Stephan et al., 2019; Vigna-Gómez et al., 2021b; Kummer et al., 2025; Shariat et al., 2025b). We thus choose the star’s semimajor axis from a log-normal distribution between [3-100] au, and a thermal distribution for the eccentricity (e.g., Shariat et al., 2025a). Note that this approach assumes zero (or very small) BH natal velocity (e.g., Naoz et al., 2016; Shariat et al., 2025b). Lastly, the mutual inclination between the inner and outer binary is chosen from an isotropic distribution (uniform in $\cos i$); the arguments of the inner and outer perihelions are chosen from a uniform distribution from $0-2\pi$.
@@
We solve the hierarchical triple-body equation of motion up to the hexadecapole level of approximation. There are two main reasons for the usefulness of this approximation beyond the octupole level. The first is that it allows us to integrate inner binaries with comparable masses. The second reason relates to the fact that in some of the systems, the ratio between the period of the outer orbit and the timescale of the eccentric Kozai-Lidov (EKL) cycles is comparable to or larger than the strength of the octupole term. It was demonstrated that in such systems, the octupole level is insufficient, and the next-level approximation, i.e., hexadecapole, allows for a more accurate description of the dynamics (e.g., Soderhjelm, 1975; Ćuk & Burns, 2004; Luo et al., 2016; Will, 2017, 2021; Tremaine, 2023; Klein & Katz, 2024). Additionally, we include the 1st post-Newtonian (1PN) precession of the inner and outer orbit, which can either suppress or excite eccentricity oscillations for either orbits (e.g., Naoz et al., 2013b, 2017; Will, 2017; Lim & Rodriguez, 2020). We neglect the 1.5PN and 2PN terms, which would slow down the calculation and only impact spin evolution and not significantly change the orbital precession. On the other hand, we add the 2.5PN terms, which induce orbital shrinking and circularization via GW emission according to (e.g., Peters & Mathews, 1963; Peters, 1964).
@@
where $\epsilon$ is the dimensionless parameter that appears as the prefactor of the octupole-level term in the hierarchical three-body Hamiltonian (e.g., Lithwick & Naoz, 2011; Naoz et al., 2013a). Although various alternative stability criteria exist in the literature (e.g., Mardling, 2010; Mushkin & Katz, 2020; Zhang et al., 2023), the $\epsilon$ criterion has been shown to be broadly consistent with many of these (e.g., Naoz & Fabrycky, 2014). Furthermore, while the inclusion of higher-order terms, such as the hexadecapole-level approximation, can extend the validity of secular dynamics to more compact systems (e.g., Will, 2017), the $\epsilon\leq 0.1$ threshold remains a conservative and robust choice.
@@ 4.3 Gaia-BH-like systems
To estimate the efficiency of this channel in producing Gaia-BH-like systems, we begin by estimating the total number of LVK sources in a galaxy, assuming that galaxies with masses comparable to the Milky-Way (MW) produce a significant fraction of the LVK rate (if they do not, the estimates below of Gaia-BH-like systems in the MW are upper limits). Given the number densities of massive galaxies $\sim 0.01-0.001$ Mpc-3 (e.g., Conselice et al., 2016), over a Hubble time, the expected number of BH mergers per galaxy is $N_{\rm BBH}\sim 3\times 10^{4-5}$. Thus, to estimate the number of Gaia-like systems, we multiply this number by $f_{\rm BH}\times f_{1}$, i.e., the fraction that our initial conditions represent from the total simulated binary BH population. Then, we multiply this by $f_{P_{\leq 10}}\sim$$0.08$ , which is the fraction of systems out of all runs that remain bound after the kick, have a period smaller than $10$ years, and have a pericenter larger than $r_{\rm circ}$. Thus, the expected Gaia-BH-candidates from this channel are
@@
The total number of Gaia-BH-like systems in our Milky Way is estimated to be about 20,000 (Nagarajan et al., 2025), thus, this channel may contribute to 1-10% of Gaia-BH-like systems. In our models, these Gaia-BH-like populations have an average separation of $\sim 7$ au and eccentricity of $0.6$. This estimate does not depend on the spin of the BHs, but it is sensitive to the alignment of the BHs, where misaligned BHs will result in even fewer post-kick bound systems. While the binary merger channel naturally results in spin alignment, in some cases, where the star is on a tight orbit, it torques the inner, BH binary orbit, which may result in misalignment (e.g. Liu & Lai, 2018). We reserve this part of the investigation for future work.

### v2

@@ 1 Introduction
Stellar-mass black holes (BHs) can be detected through multiple observational methods, each of which potentially probes different evolutionary pathways. The oldest method relies on electromagnetic (EM) signatures, most notably X-ray emission from accretion disks in X-ray binaries. These systems, discovered in the 1960s, provided the first indirect evidence for stellar-mass BHs (e.g., Bolton, 1972; Remillard & McClintock, 2006; MacLeod & Grindlay, 2023). More recently, the detection of gravitational waves (GWs) from merging binary BHs by the LIGO–Virgo–KAGRA (LVK) Collaboration has revolutionized our understanding of the BH mass distribution (e.g., Abbott et al., 2016), revealing a population of heavy BHs (up to $\sim 100$ M) without any known EM activity [^1] A handful of LVK BHs have been suggested to coincide with EM flares in AGN (Graham et al., 2023; Huang et al., 2025) but their brightness is difficult to explain (Tagawa et al., 2024) and the significance of the spatial associations are controversial (Veronesi et al., 2025).. The third avenue of detection, based on astrometric measurements of orbital motion, has been enabled by Gaia, which has opened a new observational window into detached BH binaries (El-Badry et al., 2023b). Stellar-mass BHs can also be detected by microlensing (e.g., Agol et al., 2002; Lu et al., 2016; Lam et al., 2022; Sahu et al., 2022) or by combined radial velocity and photometric obeservations of non-accreting BHs in binaries (e.g., Thompson et al., 2019; Liu et al., 2019b; Clavel et al., 2021; Chawla et al., 2024), but our focus in this paper is on the first three channels summarized above.
@@
We stress that while hierarchical triples have been proposed as a formation channel for merging binary BHs in the field, their overall contribution to the LVK-detected merger rate is typically estimated to be small and suppressed by a factor of $\sim 5$$30$ relative to the estimated detected rate (e.g., Antonini et al., 2016, 2017; Silsbee & Tremaine, 2017; Liu et al., 2019a; Dorozsmai et al., 2024; Kummer et al., 2025). Although it is important to note that each of these studies has surveyed a limited part of the parameter space. Nonetheless, triple evolution may naturally account for certain systems that are otherwise difficult to explain, such as BH mergers in the mass gap, and other puzzles (e.g., Vigna-Gómez et al., 2021; Lu et al., 2021; Martinez et al., 2022; Dorozsmai et al., 2025). In this work, however, we focus on a complementary scenario: BH binaries that merge through isolated binary evolution, with a tertiary companion that does not play a significant dynamical role in driving the merger. In particular, the tertiary star’s presence is relevant only after the inner binary merges, when the system receives a recoil kick. Our setup, therefore, explores the consequences of a passive tertiary in an otherwise binary-driven channel. We assume throughout this paper that the distant tertiary companion remains bound during the formation of the two BHs in the inner binary, and focus on the dynamics that follow when the inner BH binary merges. The validity of this assumption is sensitive, of course, to the uncertain mass ejection and natal kick during BH formation. Recent detections of X-ray binaries with distant companions imply that at least some distant companions remain bound during BH formation (e.g., Burdge et al., 2024; Shariat et al., 2025b).
@@
Figure 2: Two examples of the orbital outcome of a recoil kick on an orbit. We consider a system of two low-spin BHs with $S_{1}=S_{2}=0.01$, and spin-orbit angles of $2^{\circ}$, $1^{\circ}$, $\Omega_{s1}=90^{\circ}$ and $\Omega_{s2}=270^{\circ}$. The mutual inclination is set to $0.1^{\circ}$. Before the kick, the star has a semi-major axis of $20$ au and eccentricity $e_{\star}=0.9$ (right column), and $e_{\star}=0.83$ (left column). The argument of periapsis of the inner (outer) binary at the time of the merger is $300^{\circ}$ ($20^{\circ}$). These angles are relevant for the rotation of the various vectors to the invariant plane. See text for more details. We assume that the kick took place when the star’s phase was at $10^{\circ}$ (right column), and $5^{\circ}$ (left column). Various post-merger quantities are shown as a function of the mass ratio of the two original BHs, $q$. The bottom row shows the closest approach, Equation (3). Overplotted are the tidal radius, $r_{\rm circ}$, and the Roche radius $r_{\rm Roche}$. We expect EM signatures for BHs that approach their stellar companion within these radii. The middle row depicts the post-kick semi major axis $a_{\star,n}$, solid red lines and pericenter $a_{\star,n}(1-e_{\star,n})$, dashed purple lines. Overplotted are the initial (pre-kick) semi-major axis and pericenter of the star, solid and dashed gray lines, respectively. Lastly, Gaia-BH1’s semi-major axis and pericenter ($a_{\rm Gaia~BH1}$, and $r_{p,\rm Gaia~BH1}=a_{\rm Gaia~BH1}(1-e_{\rm Gaia~BH1})$) are also overplotted, in pink and light purple lines (El-Badry et al., 2023b). The top row shows the normalized velocity $u_{k}=v_{\rm kick}/v_{r}$, and $\sin\gamma$, where $\gamma$ is the angle between the radius vector and the post-kick velocity vector. The grid highlights that when $u_{k}=1$, $\gamma\to 0^{\circ}$, thus $R_{c}\to 0$. Note that here a lower (higher) pre-kick eccentricity, while keeping all of the other parameters constant, yields a wider, more easily unbound (tighter) post-kick binary. The opposite trend is expected for a kick that takes place at apo-center, see Equations (5) and (7).
@@ 2.1 The Triple Configuration
Throughout this work, we consider a stellar-mass BH binary, $m_{1}$ and $m_{2}$, where the mass ratio is $q\equiv m_{1}/m_{2}\leq 1$, where “1” stands for the less massive component. The total mass of the binary is $M=m_{1}+m_{2}$, semi-major axis $a_{\rm BHs}$ and eccentricity $e_{\rm BH}$, associated with angular momentum $L_{\rm BHs}$. We define the spin-orbit angle of mass $m_{j}$ as $i_{sj}=\hat{h}\cdot\hat{S}_{j}$, where $\hat{h}$ is the unit vector along the BH’s orbital angular momentum, and $j\in\{1,2\}$. The corresponding spherical coordinate angles are $\Omega_{sj}$ for each mass. Further, this binary is orbited by a star with a mass $m_{\star}=1$ M, and semi-major axis $a_{\star}$ and eccentricity $e_{\star}$. We note that while triples are observed to have a twin excess (Shariat et al., 2025a), the long lifetime of 11 M benefits the calculation. In other words, the star is less likely to evolve before the BH binary merges. A full IMF calculation for the tertiary mass is beyond the scope of this paper. The frame of reference considered here is the inertial frame, for which the $z$-axis is parallel to the total angular momentum. After the BHs merge, they typically receive a recoil kick due to the anisotropic emission of GWs.
@@ 2.2 Impact on the Stellar Orbit
We adopt the Lousto et al. (2010, 2012) fitting formulae, which provide the recoil kick velocity. See Appendix A for the relevant set of equations. In this Letter, we focus on the merger channel driven by binary stellar evolution. This channel implies that the BH spins are aligned with its angular momentum. As a result, a well-known outcome emerges: the recoil kick is directed within the plane of the BH binary (e.g., Lousto et al., 2012). Note that since natal kicks may induce misalignment, the recoil kick may be tens of degrees off-axis. We discuss the effects of the spin-alignment approximation in Section 4.3.
@@
[…]
@@ 3 A population study
As a proof of concept, we focus on a population of isolated BH binaries that will ultimately become LVK sources, and add a stellar tertiary to them. The isolated binary channel involves a common-envelope, stable mass transfer phase, or chemically homogeneous evolution (e.g., Belczynski et al., 2002, 2007; Dominik et al., 2012, 2015; Stevenson et al., 2017; de Mink & Mandel, 2016; Mandel & de Mink, 2016; Gallegos-Garcia et al., 2021). We thus adopt binary BH orbital configurations that isolated binary population synthesis work suggests will merge. Specifically, we are motivated by the angular momentum constraints from Figure 13 in Kruckow et al. (2018), which represent the orbital configurations of two BHs, right after the formation of the second BH. Specifically, we choose the initial BH binary eccentricity from a uniform distribution between 0 -1, and apply the angular momentum constraints on the binary’s semi-major axis. Additionally, the population is limited to a merger time of $\leq 10$ Gyr (which reduces the population by a factor of 2). Note that an eccentric BH binary torques the star’s inclination via the iEKL; at the time of merger, the BH binary dissipates its eccentricity via GW emission.
For simplicity, the BH masses are adopted from a uniform distribution, each between 91009-100 M. We note that the choice of mass distributions here may affect the resulting rates at the order of $\sim\pm 10-20\%$, as tested via toy models adopting a double power law, with Gaussian peaks for the LVK mass distribution, following Callister & Farr (2024) and The LIGO Scientific Collaboration et al. (2025). These tests are omitted here to avoid clutter. Given these and other uncertainties, we reserve a more detailed analysis for future endeavors.
@@
For the outer orbit, we are motivated by triple conditions that include post-main-sequence stellar evolution for massive stars (e.g., Naoz et al., 2016; Stephan et al., 2019; Vigna-Gómez et al., 2021; Kummer et al., 2025; Shariat et al., 2025b). We thus choose the star’s semimajor axis [^2] Note that the aforementioned studies yield wide tertiary orbital separations. Tighter configurations may increase the rates of the various outcomes. from a log-normal distribution between [3-100] au, and a thermal distribution for the eccentricity (e.g., Shariat et al., 2025a). Note that this approach assumes zero (or very small) BH natal velocity (e.g., Naoz et al., 2016; Shariat et al., 2025b). Lastly, the mutual inclination between the inner and outer binary is chosen from an isotropic distribution (uniform in $\cos i$); the arguments of the inner and outer perihelions are chosen from a uniform distribution from $0-2\pi$.
@@
The setup of the system is such that in most parts of the parameter space, the BH binary torques the star rather than the star influencing the dynamics of the BH binary. This can yield a preferred orientation for the star’s orbital plane (e.g., Naoz et al., 2017; Hansen & Naoz, 2020). To find the possible effect of the star’s orientation, we solve the hierarchical triple-body equation of motion up to the hexadecapole level of approximation. There are two main reasons for the usefulness of this approximation beyond the octupole level. The first is that it allows us to integrate inner binaries with comparable masses. The second reason relates to the fact that in some of the systems, the ratio between the period of the outer orbit and the timescale of the quadrupole-level eccentric Kozai-Lidov (EKL) cycles is comparable to or larger than the strength of the octupole term. It was demonstrated that in such systems, the octupole level is insufficient, and the next-level approximation, i.e., hexadecapole, allows for a more accurate description of the dynamics (e.g., Soderhjelm, 1975; Ćuk & Burns, 2004; Luo et al., 2016; Will, 2017, 2021; Tremaine, 2023; Klein & Katz, 2024). Additionally, we include the 1st post-Newtonian (1PN) precession of the inner and outer orbit, which can either suppress or excite eccentricity oscillations for either orbits (e.g., Naoz et al., 2013b, 2017; Will, 2017; Lim & Rodriguez, 2020). We neglect the 1.5PN and 2PN terms, which would slow down the calculation and only impact spin evolution and not significantly change the orbital precession. On the other hand, we add the 2.5PN terms, which induce orbital shrinking and circularization via GW emission according to (e.g., Peters & Mathews, 1963; Peters, 1964).
@@
where $\epsilon$ is the dimensionless parameter that appears as the prefactor of the octupole-level term in the hierarchical three-body Hamiltonian (e.g., Lithwick & Naoz, 2011; Naoz et al., 2013a). Although various alternative stability criteria exist in the literature (e.g., Mardling, 2010; Mushkin & Katz, 2020; Vynatheya et al., 2022; Tory et al., 2022; Zhang et al., 2023), the $\epsilon$ criterion has been shown to be broadly consistent with many of these (e.g., Naoz & Fabrycky, 2014). Furthermore, while the inclusion of higher-order terms, such as the hexadecapole-level approximation, can extend the validity of secular dynamics to more compact systems (e.g., Will, 2017), the $\epsilon\leq 0.1$ threshold remains a conservative and robust choice.
@@ 4.3 Gaia-BH-like systems
To estimate the efficiency of this channel in producing Gaia-BH-like systems, we begin by estimating the total number of LVK sources in a galaxy, assuming that galaxies with masses comparable to the Milky-Way (MW) produce a significant fraction of the LVK rate (if they do not, the estimates below of Gaia-BH-like systems in the MW are upper limits). We note that the LVK binaries are often thought to have originated from a low metallicity environment (e.g., Dominik et al., 2012; Rodriguez et al., 2015; Fishbach & Kalogera, 2021). However, the strong metallicity dependence of BH binary formation is still uncertain (e.g., van Son et al., 2025). Thus, estimating the galactic BHs from LVK mergers is done here just to provide an order of magnitude sense of the number of galactic BHs where post merger recoil could be important for setting the systems properties.
Given the number densities of massive galaxies $\sim 0.01-0.001$ Mpc-3 (e.g., Conselice et al., 2016), over a Hubble time, the expected number of BH mergers per galaxy is $N_{\rm BBH}\sim 3\times 10^{4-5}$. Thus, to estimate the number of Gaia-like systems, we multiply this number by $f_{\rm BH}\times f_{1}$, i.e., the fraction that our initial conditions represent from the total simulated binary BH population. Then, we multiply this by $f_{P_{\leq 10}}\sim$$0.08$ , which is the fraction of systems out of all runs that remain bound after the kick, have a period smaller than $10$ years, and have a pericenter larger than $r_{\rm circ}$. Thus, the expected Gaia-BH-candidates from this channel are
@@
The total number of Gaia-BH-like systems in our Milky Way is estimated to be about 20,000 (Nagarajan et al., 2025), thus, this channel may contribute to 1-10% of Gaia-BH-like systems. In our models, these Gaia-BH-like populations have an average separation of $\sim 7$ au and eccentricity of $0.6$. This estimate does not depend on the spin of the BHs, but it is sensitive to the alignment of the BHs, where misaligned BHs will result in even fewer post-kick bound systems. A small misalignment for the less (more) massive BH $\lesssim$ 10 (5) deg, has a negligible effect on the results. While the binary merger channel naturally results in spin alignment, in some cases, where the star is on a tight orbit, it torques the inner, BH binary orbit, which may result in misalignment (e.g. Liu & Lai, 2018). We reserve this part of the investigation for future work.